OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (anonymous):

y + 2 = -3(x - 7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

yes y not

OpenStudy (anonymous):

here you have the slope ans a point slope = m = -3 and point (7,-2) we have point slope form for a line is y-y1 = m(x-x1) here (x1,y1) = (7,-2) and m = -3 just substitute these vales in the above equation and get your answer

OpenStudy (anonymous):

thank u so much ure the best v

OpenStudy (anonymous):

welcomeee